# Optimal estimation for regression discontinuity design

### v1

@@ 4.2 Two-sided test and confidence interval
Next, we construct a uniformly valid two-sided test and confidence interval by using the one-sided test proposed in Section 4.1. We consider the following null and alternative hypotheses:
@@
where the critical values $\gamma_{l}$ and $\gamma_{r}$ must satisfy $P_{\bm{p}}(\hat{\tau}\not\in[\gamma_{l},\gamma_{r}])\leq\alpha$ under $H_{0}$. Hence, we need to choose the critical values satisfying
@@
where $\pi_{r}(\gamma_{r})\equiv\max_{\bm{p}\in\mathcal{P}(\tau_{0})}P_{\bm{p}}(\hat{\tau}>\gamma_{r})$ and $\pi_{l}(\gamma_{l})\equiv\max_{\bm{p}\in\mathcal{P}(\tau_{0})}P_{\bm{p}}(\hat{\tau}<\gamma_{l})$. We can calculate $\pi_{r}(\gamma_{r})$ as in Section 4.1 and $\pi_{l}(\gamma_{l})$ in a similar way. We then propose the following critical values $\gamma_{r}^{\ast}(\tau_{0})$ and $\gamma_{l}^{\ast}(\tau_{0})$:
@@
so that the critical values $\gamma_{r}^{\ast}(\tau_{0})$ and $\gamma_{l}^{\ast}(\tau_{0})$ satisfies (4.4).
We obtain the confidence region of $\tau$ by inverting the testing procedure. We define $\widehat{CR}_{1-\alpha}$ as the set of the hypothetical values that are not rejected by the proposed two-sided test, that is
@@
In other words, this confidence region is valid in finite samples uniformly over the Lipschitz class.
This confidence region is an interval. As discussed in Remark 4.1, $\gamma_{r}^{\ast}(\tau_{0})$ is increasing in $\tau_{0}$. Similarly, $\gamma_{l}^{\ast}(\tau_{0})$ is also increasing in $\tau_{0}$. Suppose that $t_{1}<t_{2}$ and $t_{1},t_{2}\in\widehat{CR}_{1-\alpha}$. Then, for any $t\in[t_{1},t_{2}]$, we obtain
@@
Hence, any $t$ within the interval $t_{1},t_{2}]$ must be contained in the confidence region $\widehat{CR}_{1-\alpha}$, which means that $\widehat{CR}_{1-\alpha}$ is an interval. Consequently, searching for the boundary points of $\widehat{CR}_{1-\alpha}$ suffices to construct the confidence interval.
@@
Using this algorithm, we can avoid calculating the critical value $\gamma^{\ast}_{r}(\tau_{0})$ for every $\tau_{0}\in[-1,1]$. We can calculate the right boundary point of $\widehat{CR}_{1-\alpha}$ in a similar way.
@@ 5.1 Monte Carlo Simulation
We demonstrate the performance of our estimator relative to existing estimators in Monte Carlo simulations. We compare our estimator (rdbinary) with three different estimators: (1) the Gaussian estimator (gauss) with homoskedastic variance $\sigma_{i}^{2}=1/4$ as in Section 3.1; (2) the Xu (2017)’s estimator (rd.mnl), which is specific for multinomial outcomes including the binary-outcome case as a special case; and (3) the Calonico et al. (2014)’s estimator (rdrobust). [^6] For rd.mnl and rdrobust, we use their default specifications with bias-corrected robust estimation and inference.
We compare their performance for three sample sizes $(N\in\{50,100,500\})$ of observations whose values of the running variable are equally spaced between 1-1 and $1$. We consider the following three different models of the conditional mean of a binary dependent variable: (1) the Lee (2008) model, which is a polynomial approximation of the conditional mean for Lee (2008)’s data and is frequently used in simulation studies for RD designs; (2) the “worst-case” model, which is the parameter value $\bm{p}$ maximizing the MSE of any linear shrinkage estimator among parameter values such that $p_{0,+}=p_{0,-}=1/2$; [^7] Note that the worst-case MSE of a linear shrinkage estimator is not necessarily attained at the parameter values of this model, since $p_{0,+}$ and $p_{0,-}$ are fixed at $1/2$. and (3) the flat model, where the conditional probability is constant at $0.5$. The three designs are illustrated in Figures $5.8$. For each model, the dependent variable takes $1$ with the probability specified as mean and otherwise takes 0.
Figure 5.6: The Lee (2008) model
@@
First, we demonstrate the point estimation properties of our estimator. Tables 5.1 and 5.2 compare the root MSE and bias for the estimation of the ATE at the cutoff, computed from $3000$ replication draws. Table 5.1 compares three different sample sizes for the Lee model. For all sample sizes, our estimator has substantially smaller MSEs than the other estimators. Furthermore, the differences shrink as the sample size grows and the MSEs are relatively similar for $N=500$. The same pattern is confirmed for different designs that have different Lipschitz constants $C$. Hence, our estimator is superior to the existing estimators in small samples, while their behaviors resemble in larger samples.
In all three designs, our estimator is superior in the MSEs relative to the other existing methods. Note that our and Gaussian estimators use $C$ as if its true values are known. Nevertheless, the margin of differences is extraordinary for an extremely small sample size as $N=50$, and our estimator exhibits a favorable property in estimating the small sample RD designs.
Table 5.1: Simulation: point estimates (Lee)
@@
Table 5.2: Simulation: point estimates N=100
worst_case Lee flat-50
@@
Second, we demonstrate the inference properties of our estimator. Tables 5.3 and 5.4 compare the average length and coverage probability of the four confidence intervals, computed from $5000$ replication draws. In Table 5.3, we demonstrate that our confidence interval has shorter lengths with guaranteed coverage relative to rd.mnl and rdrobust for different sample sizes. Unlike in the point estimation results, the differences in lengths remain similar as the sample size grows. Note that the Gaussian confidence interval happened to have shorter lengths while achieving the $95\%$ coverage for the Lee design. Nevertheless, the Gaussian confidence interval does not guarantee the coverage as the coverage falls below $95\%$ for the flat design. This behavior is consistent with the fact that the Gaussian confidence interval is designed for the missspecified model where the outcomes, and hence linear estimators, follow normal distributions. Our confidence interval is, by construction, correctly specified for the binary dependent variable. Hence, our estimator is preferred when the outcome is known to be a binary variable. We also note that the rdrobust confidence interval is based on large-sample asymptotics and is not specifically designed for binary outcomes, resulting in unsatisfactory coverage properties in all designs with small samples.
Table 5.3: Simulation Results. DGP = Lee
@@
Table 5.4: Simulation Results. N = 100
worst_case Lee flat
@@ 5.2 Application
We apply our estimator to a small-sample RD study of Brollo et al. (2013). Brollo et al. (2013) exploit a regional fiscal rule in Brazil to study the impact of an additional government fiscal transfer on the frequency of corruption in local politics. In Brazil, 40 percent of the municipal revenue is the Fundo de Participação dos Municipios (FPM) which is allocated based on the population size of municipalities. Specifically, each municipality is allocated into one of nine brackets by their population levels. The bracketing fiscal rule induces population thresholds that discontinuously alter the amount of the FPM transfers. Following Brollo et al. (2013), we reduce the nine brackets into seven thresholds because of sample selection in municipalities that recorded their primary dependent variable of corruption measures.
We chose this study for two reasons. First, their primary dependent variables are binary indicators. Specificaly, they study the impact of the fiscal rule on two measures of corruption indicators:
broad corruption, which includes irregularities that could also be interpreted as bad administration rather than as overt corruption; and narrow corruption, which only includes severe irregularities that are also more likely to be visible to voters. (Brollo et al., 2013, page. 1774)
Second, their sample sizes are relatively small. Particularly within each cutoff neighborhood, the sample size is limited to less than $400$ and mostly around $100$ to $200$. In those small samples, our estimator is expected to be superior to other estimators that are based on asymptotic approximations.
The following tables exhibit our rdbinary estimates and rdrobust estimates. [^9] The original study runs global polynomial estimations for each cutoff neighborhood as well as for the whole sample by pooling across cutoff neighborhoods. Their primary estimation is the fuzzy design, but we focus on the reduced-form sharp design estimates. Tables 5.5 and 5.6 report the pooling estimates over multiple cutoffs for the broad and narrow corruption indicators. $Crot$ is a rule-of-thumb value for the Lipschitz constant $C$, which is the largest (in absolute value) slope estimate from the binscatter estimation by binsreg package (Cattaneo et al., 2024a). In all the tables, we report the point estimates and confidence intervals for three different values of the constant $C$: the rule-of-thumb $Crot$; one half of $Crot$; and $1.5$ times $Crot$.
@@
For both indicators, our rdbinary estimates appear similar to the rdrobust estimates, which are valid for large samples. The sample size is $1,202$ for the whole pooling sample and hence is large enough for the rdrobust estimator. [^10] We do not report rd.mnl estimates because rd.mnl estimates sometimes failed to select a bandwidth in this dataset, particularly for small samples. For both methods and both outcomes, the $95\%$ confidence intervals include 0. This finding is different from the original study, which reports significant positive impacts on the frequency of corruptions. This difference highlights the importance of applying local nonparametric estimations for RD designs.
By pooling samples across multiple cutoffs, we obtain a large enough sample across different cutoffs. Nevertheless, heterogeneity across different cutoffs may be of interest as the original study explores the cutoff-specific estimates. However, only a few hundreds of observations are around each individual cutoff. For such a small sample, the asymptotic approximation may not perform well.
Tables 5.7, 5.8, and 5.9 present our rdbinary and rdrobust estimates of the impact on the broad corruption for $7$ different subsamples around each individual cutoff. See Online Appendix for qualitatively similar results for the narrow corruption indicator. For all specifications, confidence intervals for each subsample are much wider than for the pooled sample. Nevertheless, our rdbinary estimates tend to offer much shorter confidence intervals than rdrobust estimates. For example, Cutoff 3 has a sample size of $225$, which is too small for rdrobust to have any insights from its estimate. On the other hand, our rdbinary estimates offer reasonable lower bounds for the impact on the broad corruption measure, which are not far negative compared to the lower bound of the confidence interval from rdrobust.
@@
rdbinary (1.5Crot) 0.330 [-0.312, 0.936] 0.081 [-0.586, 0.721] 0.246 [-0.563, 0.963]
Table 5.9: Broad: at cutoffs 5 (N = 116), 6 (N = 73), and 7 (N = 46)
@@ 6 Conclusion
Empirical studies often attempt using RD designs in small samples. However, estimation is challenging in small samples because their desired large-sample properties may be lost. A few finite-sample minimax estimators are proposed. However, those minimax estimators require the knowledge of the variance parameter, which is usually unknown.
In this study, we provide a minimax optimal estimator for RD designs with a binary outcome variable and its inference procedure. The key idea in our estimator is the following: all features of the conditional distribution, including the conditional variance, are a known function of the conditional mean function for a binary variable. For binary outcomes, our estimator relies on a single tuning parameter, the Lipschitz constant for the bound on the first derivative. Specifically, our estimator is free from specifying the conditional variance function, which is often required for minimax optimal estimators for RD designs. Our estimator is also applicable to any bounded outcome variable. Hence, we offer a practical finite-sample minimax optimal estimator for typical outcome variables, and our estimation can be the last resort for RD studies which have relatively small effective sample sizes.
We demonstrate that the estimator is superior to the existing estimators in finite samples in numerical and simulation exercises. In a numerical exercise, we show that our estimator is $5$ to 2020% more efficient in the worst-case root mean squared errors than the existing minimax optimal estimators for extremely small samples. In simulation studies, we show that our estimator has much smaller mean squared errors than the existing methods for small enough sample sizes. Furthermore, we demonstrate that our inference procedure generates shorter confidence intervals with guaranteed coverage rates than the existing methods. In the empirical application to a small-sample RD study, we document that our estimator generates similar results with the standard large-sample procedure for large enough samples but provides much more informative results for small enough samples.
Our contribution is a critical baseline for developing estimation procedures for a binary or limited outcome variable in RD designs. Recent studies such as Noack and Rothe (2024) consider bias-aware inference for fuzzy RD designs. As mentioned in Introduction, the binary treatment status is a primary dependent variable in the first stage of fuzzy designs. Applying our result is not necessarily straightforward as the first-stage estimand appears in the denominator of the target estimand. We reserve developing further extensions and generalizations of our results for future research.
@@ Appendix A Proofs
$\frac{\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\tilde{\bm{w}},\bm{\theta})}{\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\hat{\bm{w}},\bm{\theta})}$ $\leq$ $\frac{\max_{\bm{\theta}\in\Theta_{g}}\mathrm{MSE}_{g}(\tilde{\bm{w}},\bm{\theta})}{\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\hat{\bm{w}},\bm{\theta})}\ =\ \frac{\min_{\bm{w}\in\mathcal{W}}\max_{\bm{\theta}\in\Theta_{g}}\mathrm{MSE}_{g}(\bm{w},\bm{\theta})}{\min_{\bm{w}\in\mathcal{W}}\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\bm{w},\bm{\theta})}.$
First, we derive a lower bound of $\min_{\bm{w}\in\mathcal{W}}\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\bm{w},\bm{\theta})$. From Theorem 2.1 and Lemmas 2.22.3, we obtain
$\min_{\bm{w}\in\mathcal{W}}\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\bm{w},\bm{\theta})$ $=$ $\max_{\theta_{0}\in[-1/2,0]}\mathrm{MSE}(\hat{\bm{w}},\tilde{\bm{\theta}}(\theta_{0}))\ \geq\ \mathrm{MSE}(\hat{\bm{w}},\tilde{\bm{\theta}}(0))$
@@
We consider a sufficiently large $n\in\mathbb{N}$ such that $c_{0}x-n^{-\alpha}\leq F_{n}(x)\leq c_{1}x+n^{-\alpha}$ for all $x\in[0,1]$ with $\alpha=1/3$. For any $\epsilon>0$, let $N(\epsilon)\equiv\max\{i\in\{1,\ldots,n\}:\|R_{i}\|\leq\epsilon\}$. Because $\|R_{1}\|\leq\cdots\leq\|R_{n}\|$, we have $N(\epsilon)=nF_{n}(\epsilon)$ for $\epsilon\in(0,1]$. Hence, under Assumption 3.1, we obtain
$c_{0}n\epsilon-n^{1-\alpha}\ \leq\ N(\epsilon)\ \leq\ c_{1}n\epsilon+n^{1-\alpha},\ \ \ \forall\epsilon\in(0,1].$

### v2

@@ 4.2 Two-sided test and confidence interval
Next, we construct a uniformly valid two-sided test and confidence interval using the one-sided test proposed in Section 4.1. We consider the following null and alternative hypotheses:
@@
where the critical values $\gamma_{l}$ and $\gamma_{r}$ must satisfy $P_{\bm{p}}(\hat{\tau}\not\in[\gamma_{l},\gamma_{r}])\leq\alpha$ under $H_{0}$. Hence, we must choose the critical values $\gamma_{l}^{\ast}(\tau_{0})$ and $\gamma_{r}^{\ast}(\tau_{0})$ such that
@@
where $\pi_{r}(\gamma_{r})\equiv\max_{\bm{p}\in\mathcal{P}(\tau_{0})}P_{\bm{p}}(\hat{\tau}>\gamma_{r})$ and $\pi_{l}(\gamma_{l})\equiv\max_{\bm{p}\in\mathcal{P}(\tau_{0})}P_{\bm{p}}(\hat{\tau}<\gamma_{l})$. We can calculate $\pi_{r}(\gamma_{r})$ as in Section 4.1 and $\pi_{l}(\gamma_{l})$ similarly. We then propose the following critical values $\gamma_{r}^{\ast}(\tau_{0})$ and $\gamma_{l}^{\ast}(\tau_{0})$:
@@
By construction, these critical values $\gamma_{r}^{\ast}(\tau_{0})$ and $\gamma_{l}^{\ast}(\tau_{0})$ satisfy (4.4).
We obtain the confidence region of $\tau$ by inverting the testing procedure. We define $\widehat{CR}_{1-\alpha}$ as the set of hypothesized values that are not rejected by the proposed two-sided test, that is,
@@
In other words, this confidence region is uniformly valid over the Lipschitz class.
This confidence region is an interval. As discussed in Remark 4.1, $\gamma_{r}^{\ast}(\tau_{0})$ is increasing in $\tau_{0}$. Similarly, $\gamma_{l}^{\ast}(\tau_{0})$ is increasing in $\tau_{0}$. Suppose that $t_{1}<t_{2}$ and $t_{1},t_{2}\in\widehat{CR}_{1-\alpha}$. Then, for any $t\in[t_{1},t_{2}]$, we obtain
@@
Hence, any $t$ within the interval $t_{1},t_{2}]$ must be contained in the confidence region $\widehat{CR}_{1-\alpha}$, which means that $\widehat{CR}_{1-\alpha}$ is an interval. Consequently, searching for the boundary points of $\widehat{CR}_{1-\alpha}$ is sufficient for constructing the confidence interval.
@@
Using this algorithm, we can avoid calculating the critical value $\gamma^{\ast}_{r}(\tau_{0})$ for every $\tau_{0}\in[-1,1]$. We can calculate the right boundary point of $\widehat{CR}_{1-\alpha}$ similarly.
@@ 5.1 Monte Carlo simulation
We demonstrate the performance of our estimator relative to existing estimators in Monte Carlo simulations. We compare our estimator (rdbinary) with three different estimators: (1) the Gaussian estimator (gauss) with homoskedastic variance $\sigma_{i}^{2}=1/4$ as in Section 3.1; (2) the XU20171’s estimator (rd.mnl), which is specific to multinomial outcomes, including binary outcomes as a special case; and (3) the Calonico.Cattaneo.Titiunik2014’s estimator (rdrobust). [^7] We obtained the rd.mnl package files from the author’s personal website (https://sites.google.com/view/kelixu/home). For rd.mnl and rdrobust, we use their default specifications with bias-corrected robust estimation and inference.
Figure 5.6: The Lee2008 model
@@
We compare their performance for three sample sizes $(N\in\{50,100,500\})$ of observations whose running-variable values are equally spaced between 1-1 and $1$. We consider the following three different models of the conditional mean of a binary dependent variable: (1) the Lee2008 model, which is a polynomial approximation of the conditional mean for Lee2008’s data and is frequently used in simulation studies for RD designs; (2) the “worst-case” model, which is the parameter value $\bm{p}$ maximizing the MSE of any linear shrinkage estimator among parameter values such that $p_{0,+}=p_{0,-}=1/2$; [^8] Note that the worst-case MSE of a linear shrinkage estimator is not necessarily attained at the parameter values of this model, since $p_{0,+}$ and $p_{0,-}$ are fixed at $1/2$. and (3) a flat model in which the conditional probability is constant at $0.5$. The three designs are illustrated in Figures $5.8$. For each model, the dependent variable takes $1$ with the probability specified as mean and otherwise 0.
@@
First, we demonstrate the point-estimation properties of our proposed estimator. Tables 5.1 and 5.2 compare the root MSE and bias for the estimation of the ATE at the cutoff, computed from $3000$ replication draws. Table 5.1 compares three different sample sizes under the Lee model. For all sample sizes, our estimator has substantially smaller MSEs than the other estimators. Furthermore, the differences decrease as the sample size increases, and the MSEs are relatively similar for $N=500$. We note that the same pattern is confirmed for different designs with different Lipschitz constants $C$. Hence, our estimator is superior to the existing estimators in small samples, whereas their behaviors resemble in larger samples.
@@
Table 5.1: Simulation: Point Estimates (Lee)
worst-case Lee flat-50
@@
Table 5.2: Simulation: Point Estimates N = 100
Second, we demonstrate the inference properties of our proposed method. Tables 5.3 and 5.4 compare the average length and coverage probability of the four confidence intervals, computed from $5000$ replication draws. Table 5.3 shows that our confidence interval has shorter lengths with guaranteed coverage than rd.mnl and rdrobust for different sample sizes. Unlike the point estimation results, the differences in lengths remain similar as the sample size increases. Note that the Gaussian confidence interval has slightly shorter lengths while achieving the $95\%$ coverage for the Lee design. Nevertheless, the Gaussian confidence interval does not guarantee uniform coverage across designs; for instance, it falls below $95\%$ for the flat design when $N=100$. This behavior is consistent with the fact that the Gaussian confidence interval is constructed under a missspecified model, where the outcomes, and hence the linear estimators, are assumed to follow a normal distribution. In contrast, our proposed confidence interval provides guaranteed coverage, a feature that makes it preferable. We also note that the rdrobust confidence interval is based on large-sample asymptotics and not specifically designed for binary outcomes, which results in unsatisfactory coverage properties for all designs, especially with small samples.
@@
Table 5.3: Simulation: Interval Estimates (Lee)
worst-case Lee flat-50
@@ 5.2 Application
We apply our estimator to a small-sample RD study by Brollo.Nannicini.Perotti.Tabellini2013 (brolloReplicationDataPolitical2019). Brollo.Nannicini.Perotti.Tabellini2013 exploit a regional fiscal rule in Brazil to study the impact of an additional government fiscal transfer on the frequency of corruption in local politics. In Brazil, 40 percent of municipal revenue comes from the Fundo de Participação dos Municipios (FPM), which is allocated based on the population size of municipalities. Specifically, each municipality is allocated to one of the nine brackets based on its population. The bracketing fiscal rule induces population thresholds that discontinuously alter the amount of FPM transfers. Following Brollo.Nannicini.Perotti.Tabellini2013, we focus on the first seven thresholds because of sample size limitations.
This study is particularly suited for our method for two reasons. First, their primary dependent variables are binary indicators. Specifically, they study the impact of the fiscal rule on two measures of corruption indicators:
broad corruption, which includes irregularities that could also be interpreted as bad administration rather than as overt corruption; and narrow corruption, which only includes severe irregularities that are also more likely to be visible to voters. (Brollo.Nannicini.Perotti.Tabellini2013, page 1774)
Second, the sample size is relatively small. In particular, within each cutoff neighborhood, the sample size is limited to less than $400$ and is mostly around $100$ to $200$. In these small samples, our estimator is expected to be superior to other estimators that are based on asymptotic approximations.
The following tables present our rdbinary estimates and rdrobust estimates. [^10] The original study runs global polynomial estimations for each cutoff neighborhood, as well as for the whole sample by pooling across cutoff neighborhoods. Their primary estimation is the fuzzy design, but we focus on the reduced-form sharp design estimates. Tables 5.5 and 5.6 report the pooled estimates over multiple cutoffs for the broad and narrow corruption indicators. $Crot$ denotes the rule-of-thumb value for the Lipschitz constant $C$, which is the largest (in absolute value) slope estimate from the binscatter estimation using the binsreg package (CattaneoCrumpFarrellFeng_2024). In all tables, we report the point estimates and confidence intervals for three different values of the constant $C$: CrotCrot; one-half of $Crot$; and $1.5$ times $Crot$.
@@
For both indicators, our rdbinary estimates appear to be similar to the rdrobust estimates, which are valid for large samples. The sample size is $1,202$ for the entire pooling sample and, hence, is sufficiently large for the rdrobust estimator. [^11] We do not report rd.mnl estimates because rd.mnl sometimes fails to select a bandwidth in this dataset, particularly for small samples. For both methods and outcomes, the $95\%$ confidence intervals include 0. This finding differs from that of the original study, which reports significant positive effects on the frequency of corruption. This difference highlights the importance of using local nonparametric estimations in RD designs.
By pooling samples across multiple cutoffs, we obtain a sufficiently large sample. Nevertheless, heterogeneity across different cutoffs may be of interest, as the original study explores cutoff-specific estimates. However, only a few hundred observations are available around each individual cutoff, and the asymptotic approximation may not perform well in such small samples.
Tables 5.7, 5.8, and 5.9 present our rdbinary and rdrobust estimates of the impact on the broad corruption indicator for seven different subsamples around each individual cutoff. See the Online Appendix for qualitatively similar results for the narrow corruption indicator. For all specifications, the confidence intervals for each subsample are much wider than those for the pooled sample. Nevertheless, our rdbinary method tends to yield much shorter confidence intervals than rdrobust. For example, Cutoff 3 has a sample size of $225$, which is too small for rdrobust to offer any meaningful implications from its confidence interval. By contrast, our rdbinary confidence intervals offer reasonable lower bounds for the impact on the broad corruption measure, which are not very negative compared with the lower bound of the confidence interval from rdrobust.
@@
rdbinary (1.5Crot) 0.330 [-0.313, 0.936] 0.081 [-0.586, 0.721] 0.246 [-0.562, 0.963]
Table 5.9: Broad: at cutoffs 5 (N = 116) and 6 (N = 73) and 7 (N = 46)
@@ 6 Conclusion
Empirical studies often use RD designs with small samples, where estimation and inference are particularly challenging. Methods that rely on large-sample properties may not perform well in such settings. Several finite-sample minimax estimators have been proposed. However, these estimators typically require knowledge of the variance, which is generally unavailable. Furthermore, the finite-sample validity of inference procedures based on these estimators requires the normality of the outcome variable.
In this study, we provide a minimax optimal estimator for RD designs with a binary outcome variable, together with its inference procedure. The only tuning parameter of our proposed estimator is the Lipschitz constant; our estimator does not require specification of the conditional variance function, unlike existing minimax estimators in RD designs. Our estimator is also applicable to any bounded outcome variable. Thus, we offer a practical method that serves as a last resort for RD studies with relatively small effective sample sizes.
We demonstrate that the proposed estimator is superior to existing estimators in finite samples through numerical and simulation exercises. In a numerical exercise, we show that our estimator is $5$ to 2020% more efficient in the worst-case root MSEs than a feasible version of the existing minimax optimal estimators for extremely small samples. Through simulation studies, we show that our estimator has much smaller MSEs than existing methods for sufficiently small sample sizes. Furthermore, we demonstrate that our inference procedure generates shorter confidence intervals with guaranteed coverage rates compared with existing large-sample methods. In the empirical application to a small-sample RD study, we document that our method generates similar results to the standard large-sample procedure for sufficiently large samples but provides much more informative results for small samples.
Our contribution provides a critical baseline for developing estimation procedures for binary or limited outcome variables in RD designs. Recent studies, such as noack_bias_aware_2024, have considered bias-aware inference for fuzzy RD designs. As mentioned in the Introduction, the binary treatment status is a primary dependent variable in the first stage of the fuzzy design. Applying our results is not necessarily straightforward in that context because the first-stage estimand appears in the denominator of the target parameter. We reserve further extensions and generalizations of our results for future research.
@@ Appendix A Proofs
$\frac{\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\tilde{\bm{w}},\bm{\theta})}{\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\hat{\bm{w}},\bm{\theta})}$ $\leq$ $\frac{\max_{\bm{\theta}\in\Theta_{g}}\mathrm{MSE}_{g}(\tilde{\bm{w}},\bm{\theta})}{\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\hat{\bm{w}},\bm{\theta})}\ =\ \frac{\min_{\bm{w}\in\mathcal{W}}\max_{\bm{\theta}\in\Theta_{g}}\mathrm{MSE}_{g}(\bm{w},\bm{\theta})}{\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\hat{\bm{w}},\bm{\theta})}.$
First, we derive a lower bound of $\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\hat{\bm{w}},\bm{\theta})$. Since $\hat{w}_{i}=0$ if $C\|R_{i}\|\geq 1/2$ by Lemma 2.3, we obtain
$\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\hat{\bm{w}},\bm{\theta})$ $\geq$ $\mathrm{MSE}(\hat{\bm{w}},\tilde{\bm{\theta}}(0))$
@@
We present and prove a more general version of Theorem 3.1, which covers the Hölder class (2.5) described in Remark 2.4. For a fixed pair of the smoothness constants $\gamma\in(0,1]$ and $C$, we redefine $\Theta$ and $\Theta_{g}$ as
$\Theta\equiv\{\bm{\theta}\in[-1/2,1/2]^{n+1}:|\theta_{i}-\theta_{j}|\leq C\|R_{i}-R_{j}\|^{\gamma}\ \text{for all $i$ and $j$}\}$
and
$\Theta_{g}\ \equiv\ \left\{\bm{\theta}\in\mathbb{R}^{n+1}:|\theta_{i}-\theta_{j}|\leq C\|R_{i}-R_{j}\|^{\gamma}\ \text{for all $i$ and $j$}\right\},$
respectively. Let
$\hat{\bm{w}}\in\mathrm{arg}\min_{\bm{w}\in\mathcal{W}}\max_{\bm{\theta}\in\Theta}\mathrm{MSE}(\bm{w},\bm{\theta})\ \ \text{and}\ \ \tilde{\bm{w}}\in\mathrm{arg}\min_{\bm{w}\in\mathcal{W}}\max_{\bm{\theta}\in\Theta_{g}}\mathrm{MSE}_{g}(\bm{w},\bm{\theta}).$
We generalize Assumption 3.1 and Theorem 3.1 as follows.
Assumption A.1.
  • (i)
    $0\leq\|R_{1}\|\leq\ldots\leq\|R_{n}\|\leq 1$.
  • (ii)
    There exist constants $c_{1}>c_{0}>0$ such that, for any sufficiently large $n\in\mathbb{N}$, $c_{0}x-n^{-\frac{1}{2\gamma+1}}\leq F_{n}(x)\leq c_{1}x+n^{-\frac{1}{2\gamma+1}}$ for all $x\in[0,1]$, where $F_{n}(\cdot)$ is the empirical distribution of $\|R_{i}\|$ when the sample size is $n$, that is, $F_{n}(x)\ \equiv\ \frac{1}{n}\sum_{i=1}^{n}1\{\|R_{i}\|\leq x\}$.
Theorem A.1.
Under Assumption A.1, we obtain $\max_{\bm{\theta}\in\Theta}\text{MSE}(\hat{\bm{w}},\bm{\theta})=O(n^{-\frac{2\gamma}{2\gamma+1}})$ and
In the following, we prove Theorem A.1. As a preliminary step, we note that the results of Theorem 2.1 and Lemmas 2.2 and 2.3 also hold for the Hölder class. These can be established using similar arguments, after redefining $\tilde{\bm{\theta}}(t)$ and $\mathcal{W}_{1}$ as
$\tilde{\bm{\theta}}(t)\equiv\left(\tilde{\theta}_{0}(t),\tilde{\theta}_{1}(t),\ldots,\tilde{\theta}_{n}(t)\right)^{\prime}\ \text{and}\ \tilde{\theta}_{i}(t)\equiv\min\{t+C\|R_{i}\|^{\gamma},1/2\}\ \text{for $i=0,1,\ldots,n$,}$
and $\mathcal{W}_{1}\equiv\left\{\bm{w}\in\mathcal{W}_{0}:w_{i}=0\ \text{if $C\|R_{i}\|^{\gamma}\geq 1/2$}\right\}$, respectively. Similarly, the bounds derived in Lemma 3.1 hold with the upper bound replaced as follows:
$1\ \leq\ \frac{\max_{\bm{\theta}\in\Theta}\text{MSE}(\tilde{\bm{w}},\bm{\theta})}{\max_{\bm{\theta}\in\Theta}\text{MSE}(\hat{\bm{w}},\bm{\theta})}\ \leq\ \hat{u}^{-2}\left(1+\frac{C^{2}\sum_{i=1}^{n}\hat{w}_{i}^{2}\|R_{i}\|^{2\gamma}}{\frac{1}{4}\sum_{i=1}^{n}\hat{w}_{i}^{2}}\right).$ (A.1)
Also, the optimal weights that minimize $\max_{\bm{\theta}\in\Theta_{g}}\mathrm{MSE}_{g}(\bm{w},\bm{\theta})$ over $\bm{w}\in\mathbb{R}^{n}$ for the Hölder class satisfy $\sum_{i=1}^{n}w_{i}=1$ and $w_{i}\geq 0$ for all $i$, similar to the optimal weights for the Lipschitz class. This follows from the arguments in Appendix D. Finally, the maximum MSE of the estimator $\hat{p}_{0}(\bm{w})$ with $\sum_{i=1}^{n}w_{i}=1$ over $\Theta_{g}$ is given by
$\max_{\bm{\theta}\in\Theta_{g}}\mathrm{MSE}_{g}(\bm{w},\bm{\theta})=C^{2}\left(\sum_{i=1}^{n}w_{i}\|R_{i}\|^{\gamma}\right)^{2}+\frac{1}{4}\sum_{i=1}^{n}w_{i}^{2}.$
Now, let $\alpha=\frac{1}{2\gamma+1}$. We consider a sufficiently large $n\in\mathbb{N}$ such that $c_{0}x-n^{-\alpha}\leq F_{n}(x)\leq c_{1}x+n^{-\alpha}$ for all $x\in[0,1]$. For any $\epsilon\geq 0$, let $N(\epsilon)\equiv\max\{i\in\{1,\ldots,n\}:\|R_{i}\|\leq\epsilon\}$. Because $\|R_{1}\|\leq\cdots\leq\|R_{n}\|$, we have $N(\epsilon)=nF_{n}(\epsilon)$ for $\epsilon\in[0,1]$. Hence, under Assumption A.1, we obtain
$c_{0}n\epsilon-n^{1-\alpha}\ \leq\ N(\epsilon)\ \leq\ c_{1}n\epsilon+n^{1-\alpha},\ \ \ \forall\epsilon\in[0,1].$